How many terms of GP . 3,3²,3³,....are needed to give the sum 120

Grade:11

1 Answers

Nishant Vora IIT Patna
7 years ago
120= \frac{3(3^n-1)}{(3-1)}
240= 3(3^n-1)
80= (3^n-1)
3^n = 81

n = 4